# Theory of Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Using Super-Resolution Operators as Background Error Covariance Matrices

### v3

@@ IV.3 Case of non-linear SR
$q\left(\bm{z}_{\text{LR}}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{|D\mathsf{F}|}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\exp\left\{-\frac{1}{2}\left\lVert\mathsf{F}(\bm{z}_{\text{LR}})-\mathsf{F}(\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}))\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\right\},$ (17a)
$p\left(\bm{z}_{\text{LR}}\mid\bm{x}\right)$ $=$ $\frac{|D\mathsf{F}|}{(\sqrt{2\pi b})^{n_{z}}}\exp\left\{-\frac{1}{2b}\left\lVert\mathsf{F}(\bm{z}_{\text{LR}})-\mathsf{F}(\bm{x})\right\rVert^{2}\right\},$ (17b)
where $|D\mathsf{F}|$ denotes the Jacobian determinant of $\mathsf{F}$.
This result suggests that SR operators can be interpreted as normalizing flows.Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) Normalizing flow is a deep generative model that converts Gaussian variables into non-Gaussian variables, generating new data that obey complicated probability distributions. In normalizing flows, invertible non-linear transformations are implemented as neural-network layers and are stacked to form the whole network to achieve complex transformations. Importantly, normalizing flows can be used to transform latent variables in VAEs.Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) This combined architecture is similar to the CVAE proposed in the present study (Section IV.5), where the SR model transforms $\bm{a}_{\text{LR}}$ into $\bm{a}_{\text{HR}}$ inside the encoder. This similarity implies the potential for using normalizing flows to further develop SRDA.
IV.4 Loss function for SR models
@@
So far, we have employed the SR operator $\mathsf{F}$, namely the trained SR model. This section presents a loss function for training this model. We do not know the true state of the system and only know the LR background $\bm{x}$ and the HR observations $\bm{y}$. The former $\bm{x}$ is spatially complete, while the latter $\bm{y}$ is spatially incomplete. The SR model is trained in this setup.
@@
We propose the following loss function to train SR models:
@@
where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space. The right-hand side clarifies that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies.Ha et al. (2019); Anwar, Khan, and Barnes (2020); Fukami, Fukagata, and Taira (2023) Since $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$, backpropagation can be performed; that is, the SR model $\mathsf{F}$ can be trained in a supervised manner, where the input is $\bm{x}$ and the target is $\bm{y}$. When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$, as shown in Hastie et al.Hastie, Tibshirani, and Friedman (2009) This fact suggests that the SR model can make unbiased inferences if the median of the observation error distribution is zero.
IV.5 Implementation of SRDA by CVAE
@@
The efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, similar to Section III.3. Thus, SRDA can be implemented using CVAEs.
@@
The CVAE for SRDA consists of an encoder and a decoder (Fig. 3). The difference from the CVAE in Fig. 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously. Here, we employ the SR model described in Section IV.4, which is trained with the loss function in Eq. (18). This encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$ and estimates the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$. The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ by calculating the variances. The decoder acts as the observation operator, as in Fig. 2. The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ as the loss function.
@@
Figure 3: Architecture of the CVAE for SRDA. The encoder estimates the HR analysis vector $\bm{a}_{\text{HR}}$ using the SR model. The encoder also estimates the covariance $V$ for $\bm{a}_{\text{HR}}$. The variables $\bm{a}_{\text{HR}}$ and $V$ give a sample of the true state $\bm{\hat{z}}$ in Eq. (14c). The decoder acts as the observation operator and reconstructs the observation vector $\bm{\hat{y}}=\mathsf{H}(\bm{\hat{z}})$.
@@
Strictly speaking, a different SR model should be used within the encoder because the SR model in Section IV.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$. This difference in input may introduce bias to $\bm{a}_{\text{LR}}$. In other words, the trained SR model assumes $\bm{x}$ as input and may lead to a bias of $\bm{a}_{\text{LR}}$ similar to that of $\bm{x}$. On the other hand, the output $\bm{a}_{\text{HR}}$ ($=\mathsf{F}(\bm{a}_{\text{LR}})$) is expected to be less biased because the SR-model output is unbiased if the observation error distribution has a zero median (Section IV.4). Thus, $\bm{a}_{\text{HR}}$ is used as the initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section IV.6). One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model during the training of the CVAE. For simplicity, however, the present study does not use such an advanced learning method, which remains a future research topic.
IV.6 Sequential SRDA by the CVAE
@@
We introduce a sequential SRDA method applicable to CFD simulations. The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model. The forecast and assimilation are executed alternately, as in other statistical DA methods,Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) such as 3D-Var. This assimilation is performed using a trained CVAE’s encoder (Fig. 3).
@@
Figure 4 shows a schematic of the sequential SRDA. Here, we assume a constant assimilation interval $\Delta T$. At every assimilation time step, the encoder assimilates the HR observations $\bm{y}$ into the LR forecast state $\bm{x}$, thereby estimating the HR analysis $\bm{a}_{\text{HR}}$. This output is resized to the LR using an algebraic method, such as linear interpolation. In the next assimilation cycle, the resized analysis is used as the initial condition for the LR fluid model, and forecasts are computed by numerical integration over $\Delta T$. The forecast at the final time step is fed to the encoder to infer the subsequent HR analysis.
@@
The decoder is not used in the sequential SRDA; rather, it is used only in the training phase to learn the observation operator (Section IV.5). In contrast, the trained SR model can be used in the test phase (i.e., the operational phase) to super-resolve LR forecast states. This SR operation is independent of assimilation, allowing for the generation of HR forecast states at any time step (Section VI.3).
V Methods
We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs).Fletcher (2023a) The ground truth (or the nature run) is generated by an HR fluid model, from which pseudo (or synthetic) observations are computed. The SRDA employs an LR version of the fluid model, and the resultant forecast and analysis are compared with the ground truth. A simple jet stream system was adopted for these experiments, as a proof of concept for the SRDA.
V.1 Fluid simulation
An idealized barotropic ocean jet, proposed by David et al.,David, Marshall, and Zanna (2017) was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi.Yasuda and Onishi (2023) The experimental setup is briefly explained here. See the above references for further details.
The governing equations in the periodic channel $(x,y)$ are as followsDavid, Marshall, and Zanna (2017)
@@ V.1 Fluid simulation
where $x$ is the channel-wise direction and $y$ is the transverse direction ($x\in[0,2\pi]$ and $y\in[0,\pi]$). Equation (19a) describes the evolution of vorticity $\omega$ over time $t$. The left-hand side represents the advection of $\omega$ and the beta effect,Vallis (2017a) while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$. Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively.
@@
where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero. This setup falls within the parameter regime of mixing barriers with strong eddies,David, Marshall, and Zanna (2017) where coherent vortices persist in a statistically steady state (Section VI.1).
The initial condition was a zonal jet superimposed with random perturbations. The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion,Vallis (2017b) which is a necessary condition for barotropic instability. Perturbations were added to the vorticity field for each wavenumber component, with phase and amplitude drawn from Gaussian distributions. The simulation data for deep learning were generated by varying these random perturbations.
Numerical integration was conducted using the modified Euler and pseudo-spectral methods at two spatial resolutions: LR and HR. Table 1 provides detailed configurations for the LR and HR. All simulations ran from $t=0$ to $20$ with output time intervals of $0.25$. The CFD models using the LR and HR configurations are referred to as the LR and HR fluid models, respectively.
Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR. The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule).Canuto et al. (1988)
@@ V.2 Synthetic observations
Synthetic observations were generated from the HR simulations. We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere.Lin and Yang (2020); Fletcher (2023b) Observed grid points were randomly subsampled at constant spatial intervals, namely $8\times 8$ grids, which means that 1.56% ($=1/8^{2}$) of grid points are observed. To demonstrate that the SRDA can address changes in observed points, the subsampled points were varied in time by random shifts. For the unobserved points, Not a Numbers (NaNs) were assigned and replaced with missing values in the subsequent preprocessing (Section V.3). To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity. The noise has a mean of 0 and a standard deviation of 0.1. This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
V.3 Neural networks
@@
We outline the network architectures of an SR model and a CVAE for the sequential SRDA (Section IV.6). Hyperparameters are provided in Appendix and the full implementation is available at our GitHub repository (see Data Availability Statement). PyTorch 1.11.0Paszke et al. (2019) was used for the implementation.
@@
The SR model architecture is shown by the top panel in Fig. 5. This model resembles the neural network for SR known as VDSR, proposed by Kim et al.Kim, Lee, and Lee (2016) The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR. First, this input is nonlinearly transformed through a stack of convolution and ReLU layers. Upsampling is then executed by two blocks that include pixel shuffle layers.Shi et al. (2016) Each block doubles the size of each dimension, resulting in a fourfold increase in resolution. Additionally, a skip connectionHe et al. (2016) is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output. This skip connection allows the SR model to capture small-scale patterns, that is, discrepancies between the LR input and the HR output.
@@
Figure 5: Network architectures of the SR model and the CVAE. The latter consists of the encoder and the decoder. The label “Res Block” refers to a residual block, and “Res Block x6” indicates the stacking of six Res Blocks. As in the typical implementation of VAEs,Kingma and Welling (2019) the variance output is regarded as $\ln V$ because neural networks can generally output both positive and negative values. Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp[1/2\ln V]$ provides a realization of the true state $\bm{\hat{z}}$ in Eq. (14c).
@@
The CVAE consists of an encoder and a decoder (Fig. 5). The encoder performs SRDA and integrates the inputs, an LR forecast and an HR observation, into the HR analysis. The HR observation is first downsampled through convolution to align the resolution to the LR (totally, downsampled by a factor of 4). The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$. Then, $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$. The variance for $\bm{a}_{\text{HR}}$ is denoted as $V$ in Eq. (12) and is also estimated by the encoder. Although $V$ and $\bm{a}_{\text{HR}}$ can be treated independently in the encoder (Section IV.5), the present study computes $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of neural network parameters. Moreover, we employed the feature extractor in the trained SR model to compute $V$ (Feature Extractor-1 in Fig. 5).
@@
The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$. This operation is achieved through a stack of convolutional and Leaky ReLUMaas et al. (2013) layers. We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), although this replacement slightly decreases the inference accuracy. This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section V.2); that is, the observation operator is linear. Although the proposed SRDA can theoretically address nonlinear observation operators (Section IV), such a nonlinear case is not examined in this proof of concept, which remains a topic for future research.
@@
The clipping function, ${\text{clip}}_{[0,1]}(z)=\min\{1,\max\{0,z\}\}$ ($z\in\mathbb{R}$), restricts the value range to the interval $[0,1]$. The parameters $m_{\omega}$ and $s_{\omega}$ are determined such that 99.9% of the vorticity values fall within the $[0,1]$ range. After applying Eq. (21), zero is assigned to the HR grid points without observations, and zero is regarded as a missing value. This clear interpretation of zero facilitates feature extraction from input data.Yasuda, Onishi, and Matsuda (2023)
V.4 Training and testing methods for the neural networks
V.4.1 Training method
We adopted an offline strategy in which training data are pre-generated [Fig. 6(a)]. The training data consist of pairs of HR observations and LR forecasts, as discussed in Section IV.5. The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain. After time integration over $\Delta T=1$, the resultant LR forecasts were used in training. This $\Delta T$ equals the assimilation interval, as explained below, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet. The corresponding observations were obtained by subsampling and adding Gaussian noise (Section V.2). Figure 6(a) denotes pairs of the LR forecast and the HR observations as stars. By varying the random initial perturbations (Section V.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.

### v4

@@ 4.3 Case of nonlinear SR
$l_{\text{SRDA-NL}}=\frac{1}{2r}\left\lVert H\mathsf{F}(\bm{a})-\bm{y}\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{F}(\bm{a})-\mathsf{F}(\bm{x})\right\rVert^{2},$ (16)
where SRDA-NL stands for SRDA with nonlinear SR. For simplicity, we set the observation operator as a matrix ($\mathsf{H}=H$) and denote $\bm{a}_{\text{LR}}$ as $\bm{a}$. The minimizer $\bm{a}$ can be estimated by gradient descent in an iterative manner:
$a^{(n+1)}_{i}$ $=$ $a^{(n)}_{i}-\eta^{(n)}\frac{\partial}{\partial a^{(n)}_{i}}l_{\text{SRDA-NL}},$ (17a)
$\frac{\partial}{\partial a^{(n)}_{i}}l_{\text{SRDA-NL}}$ $=$ $\frac{1}{r}\sum_{j,k,l}\left[H_{jk}\mathsf{F}_{k}(\bm{a}^{(n)})-y_{j}\right]H_{jl}\frac{\partial\mathsf{F}_{l}(\bm{a}^{(n)})}{\partial a^{(n)}_{i}}$ (17b)
$+$ $\frac{1}{b}\sum_{j}\left[\mathsf{F}_{j}(\bm{a}^{(n)})-\mathsf{F}_{j}(\bm{x})\right]\frac{\partial\mathsf{F}_{j}(\bm{a}^{(n)})}{\partial a^{(n)}_{i}},$
where a superscript, such as $(n)$, denotes the number of iterations, and $\eta^{(n)}$ is a small positive value for the $n$-th iteration.
Eqs. (17a) and (17b) are similar to backpropagation formulas using mean squared errors as a loss function [48]. The differences from the observations (i.e., errors), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagate backward through the SR neural network $\mathsf{F}$. This backpropagation updates the current solution $\bm{a}^{(n)}$ (Fig. 4). Note that all parameters of $\mathsf{F}$ remain fixed because $\mathsf{F}$ is already trained. Since trained SR neural networks generally exhibit non-locality [30, 31, 32, 33], the backpropagation updates the analysis vector $\bm{a}$ from observations over distant grid points (Fig. 4). We demonstrate that a trained SR model actually exhibits this non-locality in Section 6.2. Equation (17b) indicates that the differences from the super-resolved background state, $\mathsf{F}(\bm{a}^{(n)})-\mathsf{F}(\bm{x})$, also induce updates through backpropagation. The ELBO (16) determines the balance between the updates from observation and background values.
Figure 4: Schematic of the non-local update by backpropagation through an SR neural network (i.e., a non-linear SR model). A large error (red) is propagated backward through yellow nodes in the hidden layers and updates the LR input values.
Minimizing Eq. (16) through backpropagation may be computationally expensive, although automatic differentiation in deep learning frameworks facilitates the calculation of $\partial\mathsf{F}/\partial\bm{a}^{(n)}$ in Eq. (17b). As in Section 3.3, we propose to solve the minimization problem using CVAEs.
4.4 Loss function for SR models
@@
So far, we have used the trained SR model $\mathsf{F}$. This section introduces a loss function for training $\mathsf{F}$. We do not know the true state of the system and only know the LR background $\bm{x}$ and the HR observations $\bm{y}$, where $\bm{x}$ is spatially complete and $\bm{y}$ is spatially incomplete. The SR model is trained in this setup.
@@
Assuming that the observation operator is a projection matrix, we propose the following loss function to train SR models:
@@
where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space ($\mathsf{H}=P$). Typically, $P$ consists of interpolation and/or subsampling operations. The right-hand side shows that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies [27, 28, 29, 41]. Backpropagation is feasible because $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$. The SR model $\mathsf{F}$ is trained using $\bm{x}$ as input and $\bm{y}$ as target. When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$ [49]. This result suggests that the SR model becomes unbiased if the median of the observation error distribution is zero.
4.5 Implementation of SRDA using CVAE
@@
An efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, as in Section 3.3, as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
@@
The CVAE for SRDA consists of an encoder and a decoder (Fig. 5). The difference from the CVAE in Fig. 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously. We use the SR model in Section 4.4, which is trained with the loss function in Eq. (18). The encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$, estimating the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$. This computation of $\bm{a}_{\text{LR}}$ [$=A_{\text{LR}}(\bm{x},\bm{y})$] emulates the result of iterative backpropagation through $\mathsf{F}$ (i.e., the result of minimizing $l_{\text{SRDA-NL}}$ in Section 4.3). The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] by calculating the variance $V$ in Eq. (12). The decoder acts as the observation operator, as in Fig. 2. The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ as the loss function.
@@
Figure 5: Architecture of the CVAE for SRDA. The encoder estimates the HR analysis vector $\bm{a}_{\text{HR}}$ using the SR model. The encoder also estimates the variance $V$ for $\bm{a}_{\text{HR}}$. The variables $\bm{a}_{\text{HR}}$ and $V$ give a sample of the true state $\bm{\hat{z}}$ in Eq. (14a). The decoder acts as the observation operator and reconstructs the observation vector $\bm{\hat{y}}=\mathsf{H}(\bm{\hat{z}})$.
@@
Rigorously speaking, a different SR model should be used within the encoder because the SR model in Section 4.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$. This difference in input may introduce bias in $\bm{a}_{\text{LR}}$. In contrast, the output $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] is expected to be less biased because the SR model is unbiased if the observation error distribution has a zero median (Section 4.4). Thus, $\bm{a}_{\text{HR}}$ is used as initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section 4.6). One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model $\mathsf{F}$ during the training of the CVAE. For simplicity, this study does not use such an advanced learning method, which remains a future research topic; instead, the parameters of $\mathsf{F}$ are fixed during the CVAE training.
4.6 Sequential SRDA using the CVAE
@@
We introduce a sequential SRDA method applicable to CFD simulations. The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model. The forecast and assimilation are executed alternately, as in other statistical DA methods [1, 2]. This assimilation is performed using a trained CVAE’s encoder.
@@
Figure 6 shows a schematic of the sequential SRDA. We assume a constant assimilation interval $\Delta T$. At each assimilation time step, the encoder assimilates the HR observations $\bm{y}$ into the LR forecast state $\bm{x}$, thereby estimating the HR analysis $\bm{a}_{\text{HR}}$. This $\bm{a}_{\text{HR}}$ is resized to the LR using an algebraic method, such as linear interpolation. In the next assimilation cycle, the resized analysis is used as the initial condition for the LR fluid model, and forecasts are computed by numerical integration over $\Delta T$. The forecast at the final time step is fed to the encoder to infer the subsequent HR analysis.
@@
The decoder is not used in the sequential SRDA; rather, it is used only in the training phase to learn the observation operator. In contrast, the trained SR model can be used in the test phase (i.e., the operational phase) to super-resolve LR forecast states. This SR operation is independent of assimilation, enabling the inference of HR forecasts at any time step.
5 Methods
We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs) [50], which are commonly used in testing DA methods for the atmosphere and ocean. The ground truth (or the nature run) is generated by an HR fluid model, from which synthetic observations are computed. The SRDA employs an LR version of the fluid model, and the resultant forecast and analysis are compared with the ground truth.
A simple oceanic jet system was adopted for these experiments as a proof of concept for the SRDA. This jet exhibits the development of small-scale flow patterns and the sensitivity to initial conditions. SR can infer such small-scale patterns, while DA can suppress error growth due to the initial condition sensitivity. This jet system is suitable for assessing both aspects of SR and DA.
5.1 Fluid simulation
An idealized barotropic ocean jet, proposed by David et al. [51], was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi [35]. The experimental setup is briefly explained here. See the above references for further details.
The governing equations in the periodic channel $(x,y)$ are as follows [51]:
@@ 5.1 Fluid simulation
where $x$ is the channel-wise direction and $y$ is the transverse direction ($x\in[0,2\pi]$ and $y\in[0,\pi]$). Equation (19a) describes the evolution of vorticity $\omega$ over time $t$. The left-hand side represents the advection of $\omega$ and the beta effect [52], while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$. Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively.
@@
where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero. This setup falls within the parameter regime of mixing barriers with strong eddies [51], where coherent vortices persist in a statistically steady state (Section 6.1).
The initial condition was a zonal jet superimposed with random perturbations. The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion [53], which is a necessary condition for barotropic instability. Perturbations were added to the vorticity field for each wavenumber component, with phases and amplitudes drawn from Gaussian distributions. The data for deep learning were generated by varying these random perturbations.
Numerical integration was conducted using the modified Euler and pseudo-spectral methods at two spatial resolutions: LR and HR. Table 1 provides detailed configurations for the LR and HR simulations. All simulations ran from $t=0$ to $20$ with output time intervals of $0.25$. The CFD models using the LR and HR configurations are referred to as the LR and HR fluid models, respectively.
Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR. The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule) [54].
@@ 5.2 Synthetic observations
Synthetic observations were generated from the HR simulations. We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere. Observation points were subsampled at constant spatial intervals of $8\times 8$ grid points, which means that 1.56% ($=1/8^{2}$) of grid points were observed. To demonstrate that the SRDA can address changes in observation points, the subsampled points were varied in time by random shifts. For the unobserved points, Not a Numbers (NaNs) were assigned and later replaced with missing values in the preprocessing (Section 5.3). To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity. The noise has a mean of 0 and a standard deviation of 0.1. This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
5.3 Neural networks
@@
We outline the network architectures of an SR model and a CVAE for the sequential SRDA. Hyperparameters are provided in the Appendix, and the full implementation is available at our Zenodo repository (see Data Availability Statement). PyTorch 1.11.0 [55] was used for the implementation.
@@
The SR model architecture is shown in the top panel of Fig. 7. This model resembles the neural network for SR known as VDSR, proposed by Kim et al. [56]. The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR. First, this input is nonlinearly transformed through a stack of convolution and ReLU layers. Upsampling is then executed by two blocks that include pixel shuffle layers [57]. Each block doubles the size of each dimension, resulting in a fourfold increase in resolution. Additionally, a skip connection [58] is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output. This skip connection allows the SR model to capture the discrepancies between the LR input and the HR output.
@@
Figure 7: Network architectures of the SR model and the CVAE. The latter consists of an encoder and a decoder. The label “Res Block” refers to a residual block, and “Res Block x6” refers to the stacking of six Res Blocks. As in the typical implementation of VAEs [7], the output of variance is regarded as $\ln V$ because neural networks can generally output both positive and negative values. Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp(1/2\ln V)$ provides a realization of the true state $\bm{\hat{z}}$ in Eq. (14a). All inputs and outputs consist only of vorticity (i.e., 1 channel).
@@
The CVAE consists of an encoder and a decoder (Fig. 7). The encoder performs SRDA and integrates the inputs, an LR forecast and an HR observation, into the HR analysis. The HR observation is first downsampled through convolution to match its resolution with the LR (downsampled by a factor of 4). The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$. This $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$. The variance for $\bm{a}_{\text{HR}}$ [$V$ in Eq. (12)] is also estimated by the encoder. Although $\bm{a}_{\text{HR}}$ and $V$ can be treated independently in the encoder (Section 4.5), we compute $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of parameters, using the feature extractor in the trained SR model (i.e., Feature Extractor-1 in Fig. 7).
@@
The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$. This operation comprises convolution and Leaky ReLU [59]. We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), though this replacement slightly decreases the inference accuracy. This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section 5.2).
@@
The clipping function, ${\text{clip}}_{[0,1]}(z)=\min\{1,\max\{0,z\}\}$ ($z\in\mathbb{R}$), restricts the value range to the interval $[0,1]$. The parameters $m_{\omega}$ and $s_{\omega}$ are determined such that 99.9% of the vorticity values fall within the $[0,1]$ range. After applying Eq. (21), zero is regarded as a missing value and is assigned to the HR grid points without observations. This clear interpretation of zero facilitates feature extraction from the input data [60].
5.4 Training and testing methods for the neural networks
5.4.1 Training method
We adopted an offline strategy in which training data were pre-generated (Fig. 8a). The training data consist of pairs of HR observations and LR forecasts, as discussed in Sections 4.4 and 4.5. The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain. This filtering completely removes all high-wavenumber components using discrete Fourier transform. After time integration over $\Delta T$ ($=1$), the resultant LR forecasts were used in training. This $\Delta T$ equals the assimilation interval, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet. The corresponding observations were obtained by subsampling. Figure 8a denotes pairs of the LR forecasts and HR observations as stars. By varying the random initial perturbations (Section 5.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.